Two packages are dropped from an airplane. A parachute can increase the cross sectional area of each packages by a factor of 31. The parachute on package 1 fails to open, and the terminal speed of package 1 is 10 m/s. The parachute on package 2 opens. What is the terminal speed of package 2 in m/s? Enter a number with one digit after the decimal point.
Answer #1

Using R = ½CρAv2

vterminal2 is proportional to 1/A, so 102 = 1/(A*31)

A = 1/3100

Then for package 2nd vterminal = sqrt(1/A) = sqrt(3100) = 55.7 m/s